I would appreciate any help to the following problem #25. Determine whether each pair of lines is parallel, perpendicular, or neither.
x + y = 5
x - y = 1
Thank you.

Answer by longjonsilver(2297)

x + y = 5 --> y = -x + 5
x - y = 1 --> y = x - 1

Written in this form, the coefficient of x is the gradient... we have -1 and +1. They are not parallel.

If (grad1)*(grad2) = -1 then the lines are perpendicular.

-1*1 is -1... so the 2 lines are perpendicular.

jon.